OpenStudy (anonymous):

Maybe it's because I'm over-analyzing it but I've wasted 20 sheets of paper trying to get the right answer for this problem, but I just don't get it A shot put is released with a velocity of 12 m/s and stays in the air for 2.0s. At what angle with the horizontal was it released? What horizontal distance did it travel?

OpenStudy (anonymous):

A shot put is released with a velocity of 12 m/s and stays in the air for 2.0s. At what angle with the horizontal was it released? What horizontal distance did it travel?" For the vertical component, v_initial = 12 sin x v_final = 0 a = -9.8 m/s^2 t = 1s Vf = Vi + at 0 = 12 sin x + -9.8(1) -12 sin x = -9.8 12 sin x = 9.8 x = 54.75° << Angle to horizontal = 54.75° >> Horizontal distance = V_x_component * t d = 12 cos 54.75 * 2 << d = 13.85 m >> https://www.google.com.pk/search?q=A+shot+put+is+released+with+a+velocity+of+12+m%2Fs+and+stays+in+the+air+for+2.0s.+++At+what+angle+with+the+horizontal+was+it+released%3F+++What+horizontal+distance+did+it+travel%3F&oq=A+shot+put+is+released+with+a+velocity+of+12+m%2Fs+and+stays+in+the+air+for+2.0s.+++At+what+angle+with+the+horizontal+was+it+released%3F+++What+horizontal+distance+did+it+travel%3F&aqs=chrome..69i57j69i64l2&sourceid=chrome&espv=210&es_sm=122&ie=UTF-8